OpenStudy (anonymous):

Find inner product of the following vector: <(1-i)/2, -(1+i)/2> and <(2+i)/3,-(2i)/3>

OpenStudy (unklerhaukus):

the inner product is the sum of the products of the corresponding components \[\langle x_1,x_2,x_3\rangle\cdot\langle y_1,y_2,y_3\rangle=x_1y_1+x_2y_2+x_3y_3\]

OpenStudy (anonymous):

HAHAH but How do I deal with the imaginary part?

OpenStudy (unklerhaukus):

just remember \(i^2=-1\)

OpenStudy (unklerhaukus):

can you show me where your geting stuck..?

OpenStudy (anonymous):

sure... I am new to open study wait

OpenStudy (anonymous):

I get this answer but it is incorrect.....

OpenStudy (anonymous):

\[=\frac{ -3i^2-3i+2 }{ 6 }\]

OpenStudy (experimentx):

what's your answer?

OpenStudy (anonymous):

before your comment...

OpenStudy (unklerhaukus):

\[\left\langle \frac{1-i}2, -\frac{1+i}2\right\rangle\cdot\left\langle \frac{2+i}3,-\frac{2i}3\right\rangle=\frac{1-i}2\frac{2+i}3+\frac{1+i}2\frac{2i}3\]

OpenStudy (experimentx):

gotta go ..

OpenStudy (unklerhaukus):

\[=\frac{(1-i)(2+i)+(1+i)2i}6\]

OpenStudy (anonymous):

Thank I try again...

OpenStudy (anonymous):

by the way is -2i

OpenStudy (unklerhaukus):

both of the second terms were negative so their product is positive

OpenStudy (unklerhaukus):

the answer is pretty neat when you get to it

OpenStudy (anonymous):

so the answer is \[\frac{ i+1}{ 6 }\] Rite?

OpenStudy (unklerhaukus):

\[\Large\color{red}\checkmark\]

OpenStudy (anonymous):

How do I mark the question being solved?

OpenStudy (anonymous):

in Open study

OpenStudy (unklerhaukus):

you could type this ``` \[\Large\color{red}\checkmark\] ```

OpenStudy (anonymous):

\[\Large\color{red}\checkmark\]

OpenStudy (anonymous):

cool

OpenStudy (unklerhaukus):

if you want to know what someone typed , right click (on some \(\LaTeX\); show math as;tex commands, and you can copy past it into ``` \[...\] ``` these brackets

OpenStudy (unklerhaukus):

you might want to check this space http://openstudy.com/study#/groups/LaTeX%20Practicing!%20%3A)